KiyahNicole07:

A couple took out a 5-year $30,000 loan to pay for for their wedding. After 5 years, the loan payments they had made to the bank amounted to $38,250. The interest rate on the loan, compounded continuously, is %. If they had taken an 8-year loan instead of a 5-year loan, they would have paid approximately $ more.

Vocaloid:

the formula for continuous interest is A = P*e^(r*t) where A is the amount paid, P is the amount taken out for a loan, r is the interest rate, t is the number of years, e is a mathematical constant so you would first plug in 38,250 for the amount paid A 30,000 for the amount taken out, P 5 for the number of years to calculate the interest rate then you would go back and plug in t = 8 to figure out the amount paid for the 8-year loan then you'd subtract amt. paid for 8 year loan, minus amt. paid for 5 year loan

Vocaloid:

still there? I am happy to help with the calculations if something is still unclear

KiyahNicole07:

its very unclear

Vocaloid:

alright, let's start with the first step. A = P*e^(rt) as I said before, A is the amount paid, so A = 38,250 P is the amount taken out for a loan, so P = 30,000 the loan period is 5 years to t = 5 plugging everything in: 38250 = 30000*e^(r*5) solve for r

KiyahNicole07:

how do you solve for r?

Vocaloid:

remember your rules for isolating a variable we want "r" by itself 38250 = 30000*e^(r*5) notice how the right side is being multiplied by 30,000. we do the opposite, so we divide both sides by 30,000 to isolate r (38250/30000) = e^(r*5) now, we have e raised to the 5r power. to undo the e, we take the natural log of both sides ln(38250/30000) = 5r now, since r is being multiplied by 5, we "undo" this by dividing both sides by r. r = ln(38250/30000) / 5 = ?

KiyahNicole07:

thank you I got my answer